OpenStudy (anonymous):

need help with 2 questions, will give a medal and fan!!!!

OpenStudy (anonymous):

11. A set of test scores is normally distributed with a mean of 175 and a standard deviation of 9. Bala received a score of 166. What was her percentile score? (Points : 4) 16% 27% 34% 68%

OpenStudy (anonymous):

and the other one: 14. Carnell is finding the equation of a median-median line. He found that the equation of the line through the first and third summary points is y = 3.5x + 72. Then he found that the equation of the line that is parallel to this line and passes through the second summary point is y = 3.5x + 54. What is the equation of the median-median line?

OpenStudy (anonymous):

i dont understand the medianmedian line one at all

OpenStudy (anonymous):

@kirbykirby here is another one :p

OpenStudy (anonymous):

@I<3mybaby here

OpenStudy (i<3mybaby):

@annas @april115

OpenStudy (anonymous_user):

For you initial question I believe the answer would be, A. 16%

OpenStudy (anonymous):

oh, i thought it was 34%, could you explain how you came upon that answer?

OpenStudy (anonymous_user):

I can't really explain.

OpenStudy (anonymous):

oh ok that is alright what about the other one?

OpenStudy (anonymous_user):

Are there any choices for your second question?

OpenStudy (anonymous):

yea here they are: y = 3.5x + 66 y = 3.5x + 63 y = 3.5x + 59 y = 3.5x + 57 i thought it would be A

OpenStudy (anonymous_user):

y = 3.5x + 63

OpenStudy (anonymous):

could you explain that one?

OpenStudy (anonymous):

but that answer isnt even a choice....

OpenStudy (anonymous):

54+12=72 check.... shouldnt that be 66? also what does the check mean?

OpenStudy (anonymous):

alright ty a ton :)!!!!!!!!!!!!!!

OpenStudy (anonymous_user):

No problem.

OpenStudy (anonymous_user):

Delighted to help.